# Scale Equivariant Graph Metanetworks

### v1

@@ A.4.2 Predicting CNN Generalization from weights
For evaluating our method on the CNN Generalization task, we select CIFAR-10-GS and SVHN-GS datasets from [70], following the evaluation of prior works (Zhou et al. [81], Zhou et al. [82], Lim et al. [44], Kofinas et al. [33]). In this experiment, however, we deviate from the path of previous evaluations and split each of these datasets into two subsets. The first one contains CNNs trained with ReLU activation function and the second networks with Tanh. We do so in order to assess our method separately for each of these activation functions as they introduce different symmetries. In order to compare with previous methods, we train them on each subset and report Kendall’s correlation $\tau$ metric [30] as originally proposed by Zhou et al. [81].
@@
Hyperparameters: For the ReLU datasets we implement scale equivariant update and message functions, while for the Tanh sign equivariant. In all of our experiments we use MSE training loss. Regarding the hyperparameters we search: batch size in $\{32,64,128\}$, hidden dimension for node/edge features in $\{64,128,256\}$. We also search learning rates in $\{0.001,0.0005,0.0001\}$, weight decay in $\{0.01,0.001\}$ and dropout in $\{0,0.1,0.2\}$. Again, we experiment with using only node positional encodings or also employing edge positional encodings and apply layer normalization within each MLP.
A.4.3 Extra symmetries of the sine activation function
@@
A.5 Distribution of Scaling coefficients
@@
A.6 Deferred Proofs
A.6.1 Characterisation of symmetries
@@
A.6.2 Theoretical analysis of ScaleGMN

### v2

@@ A.4.2 Predicting CNN Generalization from weights
For evaluating our method on the CNN Generalization task, we select CIFAR-10-GS and SVHN-GS datasets from [74], following the evaluation of prior works (Zhou et al. [85], Zhou et al. [86], Lim et al. [44], Kofinas et al. [33]). The above datasets contain CNNs trained with ReLU and Tanh activation functions. In order to assess our method separately for each of these activation functions, as they introduce different symmetries, we distinguish two different types of experiments. In the first case we deviate from the path of previous evaluations and split each of these datasets into two subsets. The first one contains CNNs trained with ReLU activation function and the second networks with Tanh. In order to compare with previous methods, we train them on each subset and report Kendall’s correlation $\tau$ metric [30] as originally proposed by Zhou et al. [85].
Heterogeneous activation functions: In the second case, we evaluate ScaleGMN on the whole CIFAR-10-GS and SVHN-GS datasets. We are able to do so, as our method does not impose any limitations regarding the homogeneity of the activation functions of the input neural networks. The sole necessary modification lies in the construction of the invariant layer. Specifically, following the notation of Eq. Scale Equiv. Net, each equivariant layer $\mathsf{f}^{k}$ is now equipped with two invariant nets, $\mathsf{ScaleInv}^{k}_{relu}$ and $\mathsf{ScaleInv}^{k}_{tanh}$, to be applied on the datapoints with ReLU and Tanh activation functions respectively. The rest of the architecture remains the same.
@@
Hyperparameters: For the ReLU datasets we implement scale equivariant update and message functions, while for the Tanh sign equivariant. In all of our experiments we use MSE training loss. Regarding the hyperparameters we search: batch size in $\{32,64,128\}$, hidden dimension for node/edge features in $\{64,128,256\}$. We also search learning rates in $\{0.001,0.0005,0.0001\}$, weight decay in $\{0.01,0.001\}$, dropout in $\{0,0.1,0.2\}$ and number of GNN layers in $\{2,3,4,5\}$. Again, we experiment with using only node positional encodings or also employing edge positional encodings and apply layer normalization within each MLP. For the Tanh datasets, we also experiment with the applying canonicalization or symmetrization of the sign symmetry. Finally, for the experiments with heterogeneous activation functions, we apply the same hyperparameter search as above.
A.4.3 INR editing
To evaluate the performance of ScaleGMN on equivariant tasks we opted for the task of INR editing. Aligning with the setup of [33], we utilize the MNIST INR dataset provided by Navon et al. [54] and evaluate our method on dilating the encoded MNIST digits. Specifically, for each image $i$ of the dataset, where $i\in[N]$, we compute the ground truth dilated image using the OpenCV image processing library and denote it as $f_{i}$, where $f_{i}:\mathbb{R}^{2}\rightarrow\mathbb{R}$. Let $\bm{\theta}_{i}$ denote the INR parameters of the $i$-th image and $f_{\text{SIREN}}$ the encoded function. Then, $f_{\text{SIREN}}(x,y;\bm{\theta}_{i})$ is the output of the INR at the $(x,y)$ coordinates, when parameterized by $\bm{\theta}_{i}$. The updated INR weights are computed as: $\bm{\theta}_{i}^{\prime}=\bm{\theta}_{i}+\gamma\cdot\text{ScaleGMN}(\bm{\theta}_{i})$, where $\gamma$ a learned scalar initialized to $0.01$. Finally, we minimize the mean squared error on the function space, that is:
$\mathcal{L}=\frac{1}{N\cdot d^{2}}\sum_{i=1}^{N}\sum_{x,y}^{d}\lVert f_{\text{SIREN}}(x,y;\bm{\theta}_{i}^{\prime})-f_{i}(x,y)\rVert_{2}^{2}$ (33)
Simply put, we compare the reconstructed new INR with the dilated ground truth image. Again, we do not apply any normalizations nor feed the model samples of the encoded input image (probe features).
Hyperparameters: The GNN used for ScaleGMN is a traditional Message Passing Neural Network [24], with carefully designed update and message functions that ensure sign equivariance. In contrast to the above experiments we do not add a final readout layer, as we do not compute any final graph embedding. We optimise the following hyperparameters: batch size in $\{64,128,256\}$, hidden dimension for node/edge features in ${64,128,256}$ and number of GNN layers in $\{8,9,10,11\}$. We also search learning rates in $\{0.001,0.0005,0.0001\}$, weight decay in $\{0.01,0.001\}$ and dropout in $\{0,0.1,0.2\}$. Moreover, we experiment with using only vertex positional encodings or also employing edge positional encodings. We apply layer normalization within each MLP and use skip connections between each GNN layer. The last two steps were proven valuable to stabilise training. Finally, for each MLP within the architecture we use SiLU activation function, one hidden layer and no activation function after the last layer. The learned scalar $\gamma$, also used in previous work Zhou et al. [85], was always initialized at the same value $0.001$ and consistently led to better results. All the experiments were conducted on NVIDIA GeForce RTX 4090.
Results: As shown in Table 4, the bidirectional variant achieves better results than all the previous works on metanetworks, without incorporating any of the additional techniques described above. Notably, ScaleGMN-B surpasses the performance of the graph-based permutation equivariant baseline NG-GNN equipped with 64 probe features [33]. Interestingly, our method demonstrated much better results as we added more GNN layers. This experimental observation possibly corroborates with the findings of Theorem 5.2. Regarding our forward variant, we can see that although it is able to perform on par with the previous baselines, it struggles to match the results achieved by the bidirectional variant. This behaviour, observed exclusively in this task, is theoretically anticipated for equivariant tasks. Specifically, in the forward case, neurons of layer $\ell$ only receive information from layers $\ell^{\prime}<\ell$, which hinders the computation of meaningful embeddings for nodes of the first layers. Taking into account that the INR models are rather shallow (only 2 hidden layers of 32 neurons each), then a big proportion of the whole set of nodes is being updated with less meaningfull information.
A.4.4 Extra symmetries of the sine activation function
@@
A.5 Additional ablation studies
A.5.1 Distribution of Scaling coefficients
@@
A.6 Additional Considerations
In the following section, we explore key questions raised by the reviewers during the evaluation process. These questions touch upon interesting aspects of our methodology that merit in-depth exploration.
Complexity and performance (runtime) degradation between ScaleGMN and ScaleGMN-B In the bidirectional case, we add backward edges to our graph and introduce a backward message function to discern the two message directions. Subsequently, we concatenate the outputs of the two message functions and apply the UPD function. Consequently, the additional complexity of ScaleGMN-B is introduced solely by the extra message function, with a complexity of $O(E)$. Given that ScaleGMN has the complexity of a standard GNN model $O(V+E)$, the final complexity of ScaleGMN-B is $O(V+2E)$.
Trainable parameters: GMN vs ScaleGMN(-B)? Going from a GMN model to ScaleGMN requires adapting the MSG and UPD functions to be scale equivariant, leading to more learnable parameters, as opposed to using plain MLPs. Another design choice of ScaleGMN that introduces more parameters is using different MLPs for the I/O nodes (A.1.4).
Are there any limitations on the choice of activations of the ScaleGMN network? Importantly, our method does not impose any limitations on the choice of the activation functions. We are able to select any activation function, because these are only applied within the MLPs of the invariant modules. As discussed in Section 5, the MLPs (equipped with non-linearities) are only applied after the canonicalization/symmetrization function. In case one chose to place activations in a different computational part of ScaleGMN, this would indeed limit their options so as not to compromise scale equivariance. However, this is not the case in our method.
Canon/symm mappings for all activation functions. In general, we cannot guarantee that canonicalization/symmetrization mappings are easy to construct for any activation function, since it is currently unknown if we can provide a general characterisation of the possible symmetries that may arise. Our results extend to all positively homogeneous activations, $\sigma(\lambda x)=\lambda\sigma(x),\lambda>0$, and all odd ones, $\sigma(-x)=-\sigma(x)$. We refer the reviewer to Table 1 of Godfrey et al. [25], where we can see that LeakyReLU also falls in the first category. Regarding polynomial activations, which demonstrate non-zero scaling symmetries, one option would be: (1) norm division to canonicalise scale, and (2) sign symm/canon as discussed in this paper. The above, cover a fairly broad spectrum of common activation functions.
A.7 Deferred Proofs
A.7.1 Characterisation of symmetries
@@
A.7.2 Theoretical analysis of ScaleGMN